In a lottery, players win a large prize when they pick four digits that match, in the correct order. A smaller prize is won if only three digits are matched. What is the probability that a player wins the large prize? What is the probability that a player wins the small prize?
Answer by stanbon(75887)

In a lottery, players win a large prize when they pick four digits that match, in the correct order. A smaller prize is won if only three digits are matched. What is the probability that a player wins the large prize? What is the probability that a player wins the small prize?
----------------------------------------------------
Prob of winning large prize:
# of ways to get 4 in the correct order: 1
# of possible outcomes:
10C4 = 210 groups of 4 digits
# of orders = 4! = 24
So, # of possible outcomes = 24*210 = 5040
-----------------------------------------------
P(winning larger prize) = 1/5040
==========================================
Prob of winning smaller prize
# ways to get 3 matched: 4C3 = 4
# of possible outcomes:
10C3*3# = 720
---
P(winning smaller prize) = 4/720 = 1/180
=============================================
Cheers,
Stan H.